Gravitational force on the surface of the moon is only 1/6 as strong as gravitational force on the earth. What is the weight in Newtons of a 10 kg object on the moon and on the earth?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Mass of the object = 10 kg
Weight of the object on earth = W = m × g
∴ W = 10 × 9.8
W = 98 N
Weight of the object on moon = 1/6th the weight on the earth.
As the gravitational force on the surface of the moon is only1/6 thas strong as gravitational force on the surface of the earth.
∴ Weight of the object on moon= 98/6 = 16.3 N
Weight on earth = 98 N
Weight on moon = 16.3 N

Thanks